\newcommand{\etalchar}[1]{$^{#1}$}
\begin{thebibliography}{AKP{\etalchar{+}}92}

\bibitem[ADPV04]{MR2059617}
J\"{u}rgen Appell, Espedito De~Pascale, and Alfonso Vignoli.
\newblock {\em Nonlinear {S}pectral {T}heory}, volume~10 of {\em De Gruyter
  Series in Nonlinear Analysis and Applications}.
\newblock Walter de Gruyter \& Co., Berlin, 2004.

\bibitem[AKP{\etalchar{+}}92]{Akhmerov1992MeasuresON}
R.~R. Akhmerov, Mikhail~I. Kamenski{\u{i}}, Alexandre~S. Potapov, Alexandra
  Rodkina, and Boris~N. Sadovskii.
\newblock {\em Measures of {N}oncompactness and {C}ondensing {O}perators}.
\newblock Operator Theory: Advances and Applications. Birkh{\"a}user Verlag,
  Basel Boston Berlin, 1992.

\bibitem[AV05]{AppellVaeth:Funktionalanalysis}
J\"{u}rgen Appell and Martin V\"{a}th.
\newblock {\em Elemente der {F}unktionalanalysis}.
\newblock Vieweg, Wiesbaden, 2005.

\bibitem[Gaj05]{10.2307/24894850}
Ljiljana Gaji{\'{c}}.
\newblock On measure of non-compactness in convex metric spaces.
\newblock {\em Filomat}, 19:1--5, 2005.

\bibitem[GD03]{MR1987179}
Andrzej Granas and James Dugundji.
\newblock {\em Fixed point theory}.
\newblock Springer Monographs in Mathematics. Springer-Verlag, New York, 2003.

\bibitem[KY16]{KUNZI20162}
Hans-Peter~A. Künzi and Filiz Yıldız.
\newblock Convexity structures in t0-quasi-metric spaces.
\newblock {\em Topology and its Applications}, 200:2 -- 18, 2016.
\newblock Workshop: Aspects of Contemporary Topology V.

\bibitem[Mac73]{machado1973}
Hilton~Vieira Machado.
\newblock A characterization of convex subsets of normed spaces.
\newblock {\em Kodai Math. Sem. Rep.}, 25(3):307--320, 1973.

\bibitem[SW99]{Schaefer:TVS}
H.~H. Schaefer and M.~P. Wolff.
\newblock {\em Topological {V}ector {S}paces}, volume~3 of {\em Graduate Texts
  in Mathematics}.
\newblock Springer-Verlag, New York, 2nd edition, 1999.

\bibitem[Tak70]{takahashi1970}
Wataru Takahashi.
\newblock A convexity in metric space and nonexpansive mappings. i.
\newblock {\em Kodai Math. Sem. Rep.}, 22(2):142--149, 1970.

\bibitem[Tal77]{talman1977}
Louis~A. Talman.
\newblock Fixed points for condensing multifunctions in metric spaces with
  convex structure.
\newblock {\em Kodai Math. Sem. Rep.}, 29(1-2):62--70, 1977.

\bibitem[Wun20]{Thesis:Wunderlich}
Henning Wunderlich.
\newblock \emph{Nonlinear {S}pectral {T}heory}.
\newblock Diploma thesis, {F}ern\-{U}niversit{\"{a}}t {H}agen, {G}ermany, April
  2020.

\end{thebibliography}

\documentclass[a4paper,10pt,final]{article}

\usepackage[utf8]{inputenc}
\usepackage[english]{babel}

\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amsthm}
\usepackage{amssymb}

\usepackage{hyperref}

\newcommand{\tvs}{t.v.s.{}}
\newcommand{\lcs}{l.c.s.{}}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}

\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{lemma}[theorem]{Lemma}

\theoremstyle{definition}

\title{Characterization of Fr\'{e}chet Spaces and Application to Hausdorff MNC}
\author{Henning Wunderlich\footnote{Dr. Henning Wunderlich, Frankfurt, Germany. E-Mail: \textit{HenningWunderlich@t-online.de}}}

\begin{document}

\maketitle

\begin{abstract}
In this short note, we give a characterization of Fr\'{e}chet spaces via properties of their metric. This allows us to prove that the Hausdorff measure of noncompactness (MNC), defined over Fr\'{e}chet spaces, is indeed an MNC. As first applications, we lift well-known fixed-point theorems for contractive and condensing operators to the setting of Fr\'{e}chet spaces.  
\end{abstract}

\section{Introduction}
In this short note, we give a characterization of Fr\'{e}chet spaces via properties of their metric. This allows us to prove that the Hausdorff measure of noncompactness (MNC), defined over Fr\'{e}chet spaces, is indeed an MNC, which was an open problem before. Recall that the defining property of an MNC is invariance under convex hulls. While the definition of the Hausdorff MNC over Fr\'{e}chet spaces is well-known, see e.g., \cite{Akhmerov1992MeasuresON,AppellVaeth:Funktionalanalysis}, a formal proof that it is actually an MNC was only known in the setting of Banach spaces, see e.g., \cite{MR2059617}. Past research on MNC in the setting of metric spaces \cite{talman1977,10.2307/24894850} required the existence of certain convex structures, based on the work of Takahashi \cite{takahashi1970}. Their existence in the context of Fr\'{e}chet spaces was not proven. We prove their existence with the help of the mentioned characterization.  

We think that our main result, the invariance of the Hausdorff MNC under convex hulls in Fr\'{e}chet spaces, will find many applications in the future. As first applications, we lift well-known fixed-point theorems for contractive and condensing operators of Darbo and Sadovski\u{\i} type to the setting of Fr\'{e}chet spaces. 

Substantial parts of this work have been taken from a submitted thesis of the author \cite{Thesis:Wunderlich}. Nevertheless, all relevant contents is either presented here or can be found in accessible references.

\section{Characterization of Fr\'{e}chet Spaces}
Recall that a \emph{metric vector space}\index{Metric Vector Space} $(E, d)$ is a vector space $E$ and a metric space $(E, d)$, equipped with a \emph{translation-invariant}\index{Metric!Translation-Invariant} metric $d$, i.e., $d(x, y) = d(x + z, y + z)$ for all $x, y, z \in E$. We need the following folklore result on such metrics.

\begin{lemma}[Folklore]\label{Lemma:AdditiveD}
Let $d$ be a translation-invariant metric. Then for all $x_{1}, x_{2}, y_{1}, y_{2} \in E$, we have
\begin{equation}
d( x_{1} + x_{2} , y_{1} + y_{2} ) \leq d( x_{1} , y_{1} ) + d( x_{2} , y_{2} )
\quad.
\end{equation}
\end{lemma}

\begin{proof}
First of all, for all $z_{1}, z_{2} \in E$, we have
\begin{equation}
d(0, z_{1} + z_{2}) \leq d(0, z_{1}) + d(z_{1}, z_{1} + z_{2}) = d(0, z_{1}) + d(0, z_{2})
\quad.
\end{equation}
Then
\begin{align*}
d( x_{1} + x_{2} , y_{1} + y_{2} )
&=
d( 0 , (y_{1} + y_{2}) - (x_{1} + x_{2}) )
=
d( 0 , (y_{1} - x_{1}) + (y_{2} - x_{2}) )
\\
&\leq
d( 0 , y_{1} - x_{1} )
+
d( 0 , y_{2} - x_{2} )
= d( x_{1} , y_{1} ) + d( x_{2} , y_{2} )
\quad.
\end{align*}
\end{proof}

The translation-invariant metric $d$ of a metric vector space $(E, d)$ induces a uniform topology on $E$, which makes $E$ a \tvs. A \tvs\ $E$ is called \emph{metrizable}\index{Space!Metrizable}, if there exists a translation-invariant metric $d$ on $E$ inducing the topology of $E$.

Recall that a \emph{Fr\'{e}chet space}\index{Fr\'{e}chet Space}\index{Space!Fr\'{e}chet} is a complete and metrizable \lcs. The differentiating property between complete and metrizable \tvs\ and \lcs\ is exactly the following.

\begin{theorem}[Characterization Fr\'{e}chet]\label{Theorem:CharFrechet}
Let $E$ be a complete and metrizable \tvs. Then $E$ is a Fr\'{e}chet space iff there exists a translation-invariant metric $d$ on $E$ such that for all $x, y \in E$ and $\lambda \in [0, 1]$ we have
\begin{equation}\label{Eq:DStrong}
d( \lambda \cdot x, \lambda \cdot y) \leq \lambda \cdot d(x, y)
\quad.
\end{equation}
\end{theorem}

\begin{proof}
We modify the proof in \cite[I.6.1]{Schaefer:TVS}. There, a pseudonorm $|x|$ is constructed by a base of $0$-neighborhoods $V_{n}$. The metric is then obtained via $d(x, y) = |y - x|$ and vice versa. As $E$ is an \lcs, we can assume that these $V_{n}$ are not only circled but absolutely-convex, and that $2 \cdot V_{n + 1} = V_{n}$. We prove $| 2^{-k} \cdot x| \leq 2^{-k} \cdot |x|$ for arbitrary $x \in E$ and $k \geq 1$. Then by dyadic expansion and the triangle inequality, we obtain $| \lambda \cdot x | \leq \lambda \cdot |x|$ for all real $\lambda \in [0, 1]$. Set $V_{H} := \sum_{n \in H}V_{n}$ for finite $H \subseteq \mathbb{N}$. Then $V_{k + H} = \sum_{n \in H} V_{k + n} = \sum_{n \in H} 2^{k} \cdot V_{n} = 2^{k} \cdot (\sum_{n \in H}V_{n}) = 2^{k} \cdot V_{H}$. Hence, $2^{-k} \cdot x \in V_{H}$ iff $x \in 2^{k} \cdot V_{H}$ iff $x \in V_{k + H}$. For the numbers $p_{H} := \sum_{n \in H}2^{-n}$ we get $p_{k + H} = 2^{-k} \cdot p_{H}$.  

Given arbitrary $\epsilon > 0$, let $H$ be such that $|x| \leq p_{H} - \epsilon$. Then $2^{-k} \cdot x \in V_{H}$ implies $x \in V_{k + H}$, and hence $| 2^{-k} \cdot x | \leq p_{k + H} = 2^{-k} \cdot p_{H} \leq 2^{-k} \cdot ( |x| - \epsilon)$. 
\end{proof}

The Lebesgue spaces $\mathcal{L}^{p}$ give nice examples to show, when this stronger inequality (\ref{Eq:DStrong}) holds and when it does not. Let $\lambda \in [0, 1]$. For $1 \leq p \leq \infty$, space $\mathcal{L}^{p}$ is a normed and thus a Fr\'{e}chet space, and we have $d(\lambda \cdot x, \lambda \cdot y) := \| \lambda \cdot (y - x)\|_{p} = \lambda \cdot d(x, y)$. In contrast, for $0 < p < 1$, space $\mathcal{L}^{p}$ is only a complete and metrizable \tvs, and not an \lcs. Here, we have $d(\lambda \cdot x, \lambda \cdot y) = \int | \lambda \cdot (y - x) |^{p} = \lambda^{p} \cdot d(x, y) > \lambda \cdot d(x, y)$ for $\lambda \in ]0, 1[$. 

\section{Application to Hausdorff MNC}
An important part of Functional Analysis is concerned with measures of noncompactness, see \cite{Akhmerov1992MeasuresON} for a systematic exposition of this topic. A measure of noncompactness quantifies the deviation of a bounded subset of a space from being compact. We remark that this notion does not make sense in Montel spaces, where bounded and compact sets are not distinct.

The most general definition is as follows, see also \cite[1.2.1]{Akhmerov1992MeasuresON}: Let $E$ be a \lcs, and let $(Q, \leq)$ be a partially-ordered set. A map $\chi \colon 2^{E} \to Q$ is called a \emph{measure of noncompactness (MNC)}\index{Measure of Noncompactness}, if for all subsets $A \subseteq E$ we have $\chi(A) = \chi(\overline{\mathrm{co}}(A))$. We note that going beyond \lcs\ to general \tvs\ does not make sense, because only for \lcs\ it is ensured that the convex hull of a compact set stays compact, see \cite[II.4.3]{Schaefer:TVS}. 

The Hausdorff MNC, $\alpha$, is a typical example. Another example, not treated here, is the Kuratowksi MNC, which is actually equivalent to the Hausdorff MNC, see \cite[1.1.1, 1.1.7]{Akhmerov1992MeasuresON}. 

Let $E$ be a Fr\'{e}chet space, and let $M \subseteq E$. The \emph{(Hausdorff) measure of noncompactness of $M$}\index{Measure of Noncompactness!Hausdorff} is defined by
\begin{equation}   
\alpha(M) := \inf \left\{ \epsilon > 0 \mid \textnormal{$M$ has a finite $\epsilon$-net in $E$} \right\}
\quad.
\end{equation}

For some function spaces, explicit formulas are known to compute the Hausdorff MNC, see e.g., \cite[1.1.9--1.1.13]{Akhmerov1992MeasuresON} or \cite[3.6--3.9]{AppellVaeth:Funktionalanalysis}.

The Hausdorff MNC, defined over Banach spaces, has the following properties, see e.g., \cite[Chapter 1, Proposition 1.1]{MR2059617}. With literally the same proofs, one can easily show that the properties also hold for the Hausdorff MNC $\alpha$, defined over a Fr\'{e}chet space $E$.

\begin{proposition}
Let $E$ be a Fr\'{e}chet space. For sets $M, N \subseteq E$, $z \in E$, and $\lambda \in \mathbb{K}$, we have
\begin{enumerate}
\item $\alpha(M) \leq \alpha(N)$ for $M \subseteq N$.
\item $\alpha(\overline{M}) = \alpha(M)$.
\item $\alpha(z + M) = \alpha(M)$, i.e., $\alpha$ is translation-invariant.
\item $\alpha(\lambda \cdot M) = |\lambda| \cdot \alpha(M)$, i.e., $\alpha$ is homogeneuous.
\item $\alpha(M) = 0$ iff $M$ is precompact.
\item $|\alpha(M) - \alpha(N)| \leq \alpha(M + N) \leq \alpha(M) + \alpha(N)$. The first inequality only holds in case both subsets are nonempty.
\item $\alpha(M \cup N) = \max\{ \alpha(M), \alpha(N) \}$.
\item $\alpha(B(z, 1)) = 1$, if $E$ is infinite-dimensional, and zero otherwise.
\item If $M_{1} \supseteq M_{2} \supseteq \ldots$ is a decreasing sequence of closed sets in $E$ with $\alpha(M_{n}) \rightarrow 0$ for $n \rightarrow \infty$, then the intersection $M_{\infty} := \bigcap_{n}M_{n}$ is nonempty and compact.
\end{enumerate}
\end{proposition}

Interestingly, this does not hold for the defining property of being an MNC. To the best of our knowledge, it seems to have been an open problem for a long time. The importance of this defining property has been stressed explicitly in \cite{Akhmerov1992MeasuresON}, see remark above Theorem 1.1.5 there.

\begin{theorem}\label{Theorem:MNC}
The Hausdorff MNC $\alpha$, defined over a Fr\'{e}chet space $E$, is indeed an MNC, i.e., we have $\alpha(\mathrm{co}(M)) = \alpha(M)$ for all $M \subseteq E$.
\end{theorem}

We give a direct proof first, and then discuss the existing literature.

\begin{proof}
As $M \subseteq \mathrm{co}(M)$, then $\alpha(M) \leq \alpha(\mathrm{co}(M))$ by item (i) from above proposition. For the other direction, let $N$ be a finite $\eta$-net for $M$, $\eta > 0$. Define $C := \overline{\mathrm{co}}(N)$. We have $d(x, z) \leq \eta$ for all $x \in \mathrm{co}(M)$ and $z \in C$. This can be seen as follows. Point $z$ is a convex combination $z = \sum_{i} \lambda_{i} \cdot z_{i}$ with $z_{i} \in N$, $\lambda_{i} \in [0, 1]$, and $\sum_{i} \lambda_{i} = 1$. Now, the subtle issue comes: Making use of Lemma \ref{Lemma:AdditiveD} in the second and Theorem \ref{Theorem:CharFrechet} in the third inequality, we have
\begin{align*}
d(x, z)
&=
d\left( (\sum_{i} \lambda_{i}) \cdot x, \sum_{i} \lambda_{i} \cdot z_{i} \right)
\leq
\sum_{i} d\left( \lambda_{i} \cdot x, \lambda_{i} \cdot z_{i} \right)
\\
&\leq
\sum_{i} \lambda_{i} \cdot d( x, z_{i} )
\leq
\sum_{i} \lambda_{i} \cdot \eta
=
(\sum_{i} \lambda_{i}) \cdot \eta
=
1 \cdot \eta
=
\eta
\quad.  
\end{align*}
In addition, set $C$ is compact, because it is a closed and bounded set in a finite-dimensional space $\mathrm{span}(N)$. As $C$ is compact, for every $\epsilon > 0$, there exists a finite $\epsilon$-net $K$ for $C$. Then $K$ is a finite $(\eta + \epsilon)$-net for $\mathrm{co}(M)$. 
\end{proof}

Concerning MNCs in metric spaces, the book \cite[Section 1.8.1]{Akhmerov1992MeasuresON} refers to a paper of Talman \cite{talman1977}, where the Hausdorff MNC is shown to be an MNC, if the underlying metric space has some special convex structure.

Recall that a \emph{Takahashi convex structure (TCS)}\index{Takahashi Convex Structure} on a metric space $(E, d)$ is a mapping $W \colon E \times E \times [0, 1] \to E$ such that
\begin{equation*}
d( u, W(x, y, t) ) \leq t \cdot d(u, x) + (1 - t) \cdot d(u, y)   
\end{equation*}
for all $u, x, y \in E$ and $t \in [0, 1]$. For its properties, see e.g., \cite{takahashi1970,machado1973,KUNZI20162}. A metric space, together with such a TCS, is then called a \emph{convex metric space}\index{Convex Metric Space}.

Talman refined this notion by going up one dimension. Let us call a \emph{Talman convex structure (T${}_{m}$CS)}\index{Talman Convex Structure} on a metric space $(E, d)$ a mapping $K \colon E \times E \times E \times I \to E$, where $I := \{ (\lambda_{1}, \lambda_{2}, \lambda_{3}) \in [0, 1] \mid \lambda_{1} + \lambda_{2} + \lambda_{3} = 1\}$, such that
\begin{equation*}
d( u, K(x, y, z, t_{1}, t_{2}, t_{3}) ) \leq t_{1} \cdot d(u, x) + t_{2} \cdot d(u, y)+ t_{3} \cdot d(u, z)   
\end{equation*}
for all $u, x, y, z \in E$ and $(t_{1}, t_{2}, t_{3}) \in I$. If the point $K(x, y, z, t_{1}, t_{2}, t_{3})$, satisfying above relationship, is uniquely determined, then $K$ is called a \emph{strong convex structure (SCS)}\index{Strong Convex Structure}. A metric space, together with an SCS, is then called \emph{strongly convex}\index{Strongly Convex Space}. 

Define the following mappings
\begin{align}
\tilde{W}(x, y, t)
&:=
t \cdot x + (1 - t) \cdot y
\quad,
\\
\tilde{K}(x, y, z, t_{1}, t_{2}, t_{3})
&:=
t_{1} \cdot x + t_{2} \cdot y + t_{3} \cdot z
\quad.
\end{align}

\begin{proposition}
Let $E$ be a Fr\'{e}chet space, with given metric $d$, which is trans\-lation-invariant and has property (\ref{Eq:DStrong}). Then mapping $\tilde{W}$ is a TCS, and mapping $\tilde{K}$ is a T${}_{m}$CS.
\end{proposition}

\begin{proof}
We have
\begin{align*}
d(u, \tilde{K}(x, y, z, t_{1}, t_{2}, t_{3}) )
&=
d(u, t_{1} \cdot x + t_{2} \cdot y + t_{3} \cdot z)
\\
&=
d(t_{1} \cdot u + t_{2} \cdot u + t_{3} \cdot u, t_{1} \cdot x + t_{2} \cdot y + t_{3} \cdot z)
\\
&\leq
d(t_{1} \cdot u, t_{1} \cdot x)
+
d(t_{2} \cdot u, t_{2} \cdot y)
+
d(t_{3} \cdot u, t_{3} \cdot z)
\\
&\leq
t_{1} \cdot d(u, x)
+
t_{2} \cdot d(u, y)
+
t_{3} \cdot d(u, z)
\quad.
\end{align*}
The proof for $\tilde{W}$ is analogous.
\end{proof}

We thus call $\tilde{W}$ and $\tilde{K}$ the \emph{obvious} convex structures for Fr\'{e}chet spaces. Interestingly, we remark that their existence was NOT obvious in the past. Takahashi \cite[Section 2, p.142]{takahashi1970} stated, without giving a reference or giving an example: ``But a Fr\'{e}chet space is not necessar[il]y a convex metric space.'' The insight of our work may be that one can always find the ``right'' metric (translation-invariant and property (\ref{Eq:DStrong})) such that a TCS exists (the obvious one).

Given a metric space $(E, d)$ and a TCS $W$, a subset $C \subseteq E$ is called \emph{$W$-convex}, iff $W(x, y, t) \in C$ for all $x ,y \in C$ and $t \in [0, 1]$. A $W$-convex subset $C \subseteq E$ is called \emph{stable}\index{Stable Set}, if $C_{r} := \{ x \in E \mid d(x, C) < r \}$ is also $W$-convex for every $r > 0$. An SCS on $E$ is \emph{stable}\index{Stable SCS}, if the set $\{ W(x, y, t) \mid t \in [0, 1] \}$ is stable for every pair $x, y \in E$. For a strongly convex metric space $E$, an SCS $W$ is stable iff every $W$-convex subset of $E$ is stable, \cite[Theorem 3.3]{talman1977}. 

We observe that for a Fr\'{e}chet space $E$ and its obvious TCS $\tilde{W}$, a subset is convex iff it is $\tilde{W}$-convex. Furthermore, every convex subset is stable. 

\begin{proposition}
Let $E$ be a Fr\'{e}chet space, with given metric $d$, which is trans\-lation-invariant and has property (\ref{Eq:DStrong}). Then every $\tilde{W}$-convex subset of $E$ is stable.
\end{proposition}

\begin{proof}
Let $C \subseteq E$ be convex, and let $r > 0$. For $x, y \in C_{r}$, there exist $\tilde{x}, \tilde{y} \in C$ such that $d(x, \tilde{x}), d(y, \tilde{y}) < r$. For $t \in [0, 1]$, define $z_{t} := t \cdot x + (1 - t) \cdot y$ and $z_{t} := t \cdot \tilde{x} + (1 - t) \cdot \tilde{y}$, respectively. As $C$ is convex, $\tilde{z}_{t} \in C$. We have
\begin{align*}
d(z_{t}, \tilde{z}_{t})
&=
d( t \cdot x + (1 - t) \cdot y, t \cdot \tilde{x} + (1 - t) \cdot \tilde{y} )
\\
&\leq
d( t \cdot x, t \cdot \tilde{x} ) + d( (1 - t) \cdot y,  (1 - t) \cdot \tilde{y} )
\\
&\leq
t \cdot d(x, \tilde{x}) + (1 - t) \cdot d(y, \tilde{y})
<
t \cdot r + (1 - t) \cdot r = r
\quad.
\end{align*}
Hence, $C_{r}$ is convex.
\end{proof}

Talman \cite[Theorem 3.7]{talman1977} proved the following

\begin{theorem}
Let $E$ be a strongly convex metric space with stable SCS. Then for all bounded subsets $M \subseteq E$, we have $\alpha(M) = \alpha(\mathrm{co}(M))$.
\end{theorem}

The only missing piece, which prevents us from applying this theorem on a Fr\'{e}chet space $E$, is the unclear uniqueness of the obvious TCS $\tilde{K}$, i.e., is $\tilde{K}$ even an SCS? We have doubts that such uniqueness holds in arbitrary Fr\'{e}chet spaces. Hence, \cite[Section 1.8.1]{Akhmerov1992MeasuresON}, just referring to Talman's paper \cite{talman1977}, does not give a conclusive answer for Fr\'{e}chet spaces.

Matters are different with the paper \cite{10.2307/24894850} of Gaji\'{c}. For a convex metric space $(E, d, W)$, she defines two properties (P) and (Q). Recall that $(E, d, W)$ has \emph{property (P)}\index{Property (P)}, if for all $x_{1}, x_{2}, y_{1}, y_{2} \in E$ and $t \in [0, 1]$, we have
\begin{equation*}
d( W(x_{1}, y_{1}, t), W(x_{2}, y_{2}, t) ) \leq t \cdot d(x_{1}, y_{1}) + (1 - t) \cdot d(x_{2}, y_{2})
\quad.
\end{equation*}
It has \emph{property (Q)}\index{Property (Q)}, if for every finite subset $F \subseteq E$, the $W$-convex hull of $F$, $W$-$\mathrm{co}(F)$, is compact.

The following is Theorem 1 in \cite{10.2307/24894850}.

\begin{theorem}[Gaji\'{c}]\label{Theorem:Gajic}
Let $(E, d, W)$ be a convex metric space with TCS $W$, satisfying properties (P) and (Q). Then for all bounded subsets $M \subseteq E$, we have $\alpha(M) = \alpha(W\textnormal{-}\mathrm{co}(M))$.
\end{theorem}
    
A Fr\'{e}chet space $E$, together with translation-invariant metric $d$ having property (\ref{Eq:DStrong}) and the obvious TCS $\tilde{W}$, clearly has properties (P) and (Q). The latter holds, because convex and $\tilde{W}$-convex sets equal, and compact sets are closed under convex hulls in every \lcs. Hence, Theorem \ref{Theorem:Gajic} gives another proof of Theorem \ref{Theorem:MNC}.

\section{Fixed-Point Theorems}
The property of being invariant under convex hulls is a crucial component in many proofs, involving the Hausdorff MNC. As first applications of our result, we prove fixed-point theorems for contractive and condensing operators of Darbo and Sadovski\u{\i} type. For more information on Fixed-Point Theory, we refer the reader to the opus magnum of Granas and Dugundji \cite{MR1987179}.

Let $E$ and $F$ be Fr\'{e}chet spaces, and let $F \colon E \to F$ be a continuous operator. Analogously to the Banach-space setting, the \emph{upper characteristic of noncompactness}\index{Characteristic!Noncompactness} is defined by
\begin{align}
[F]_{A} &:= \inf \left\{ \gamma > 0 \mid \alpha(F(M)) \leq \gamma\cdot \alpha(M), \textnormal{$M$ bounded} \right\}
\quad.
\end{align}

Operator $F$ is called \emph{$\alpha$-contractive}, if $[F]_{A} < 1$. It is called \emph{condensing}, if $\alpha(F(M)) < \alpha(M)$ for every bounded subset $M \subseteq E$ with $\alpha(M) > 0$.

\begin{theorem}[Darbo]
Let $E$ be a Fr\'{e}chet space, and let $F \colon E \to E$ be $\alpha$-contractive. Let $M \subseteq E$ be a nonempty, convex, closed, and bounded set such that $F(M) \subseteq M$. Then $F$ has a fixed-point in $M$, i.e., there exists $x \in M$ with $F(x) \in M$.
\end{theorem}

Slightly modify the proof in \cite[Section 2.3, Theorem 2.1]{MR2059617}, using Theorem \ref{Theorem:MNC} in
\begin{equation*}
\alpha( M_{ n + 1 } ) = \alpha( \overline{\mathrm{co}} F(M_{n}) ) = \alpha( \mathrm{co} F(M_{n}) ) = \alpha( F(M_{n}) ) \leq \cdots
\quad,
\end{equation*}
and applying the Theorem of Schauder-Tychonoff instead of the Theorem of Schauder \cite[II \S7.1, Theorem 1.13]{MR1987179}.

\begin{theorem}[Sadovski\u{\i}]
Let $E$ be a Fr\'{e}chet space, and let $F \colon E \to E$ be condensing. Let $M \subseteq E$ be a nonempty, convex, closed, and bounded set such that $F(M) \subseteq M$. Then $F$ has a fixed-point in $M$.
\end{theorem}

As before, slightly modify the proof in \cite[Section 2.3, Theorem 2.3]{MR2059617}, using Theorem \ref{Theorem:MNC} in
\begin{equation*}
\alpha(F(\hat{M})) = \alpha( \overline{\mathrm{co}} F(\hat{M}) ) = \alpha( \Phi(\hat{M})) = \alpha(\hat{M}) 
\quad,
\end{equation*}
and applying the Theorem of Schauder-Tychonoff instead of the Theorem of Schauder \cite[II \S7.1, Theorem 1.13]{MR1987179}.

Even more general, Gaji\'{c} \cite[Theorem 2]{10.2307/24894850} proved a fixed-point theorem for \emph{condensing}\index{Condensing} set-valued mappings $F \colon E \to 2^{E}$, i.e., $\alpha(M) > 0$ implies $\alpha(F(M)) < \alpha(M)$ for all bounded sets $M \subseteq E$.

\begin{theorem}[Gaji\'{c}]
Let $(E, d, W)$ be a complete TCS with continuous structure $W$ and satisfying properties (P) and (Q), respectively. Let $F \colon E \to 2^{E}$ be a set-valued, condensing mapping with $W$-convex values, closed graph, and bounded range. Then $F$ has a fixed point, i.e., there exists $x \in E$ such that $x \in F(x)$.
\end{theorem}

From this, using the properties of the obvious TCS $\tilde{W}$ on a Fr\'{e}chet space $E$, we immediately obtain

\begin{corollary}
Let $E$ be a Fr\'{e}chet space. Let $F \colon E \to 2^{E}$ be a set-valued, condensing mapping with convex values, closed graph, and bounded range. Then $F$ has a fixed point.
\end{corollary}
 
We think that this is just the beginning of lifting known results from Banach to Fr\'{e}chet spaces. As an outlook, we give two examples of possible future generalizations: First of all, establishing a Nussbaum-Sadovski\u{\i} degree and its properties \cite[Section 3.5]{MR2059617} for Fr\'{e}chet spaces, secondly, lifting the FMV and Feng spectra \cite[Chapters 6 and 7]{MR2059617} of Nonlinear Spectral Theory. In the latter case, the defining property of an MNC helps in establishing the closed- and boundedness of these spectra, see e.g., the proofs of Lemma 6.2 and Theorem 7.1 in \cite{MR2059617}. 

\section*{Acknowledgements}
First of all, we would like to sincerely thank Prof. Dr. Delio Mugnolo (Fern\-Universit\"{a}t Hagen) for his valuable advice and support. We would also like to encourage the reader to give us feedback. Any help is appreciated very much!

\bibliographystyle{alpha}
\bibliography{Frechet-MNC}

\end{document}